# Spacetime Quasicrystals

### v1

@@ V.1.2 Hierarchy Puzzles: Cartoon Picture
Now let us see how this picture might be helpful in explaining some the large hierarchies between three fundamental dimensional scales in nature: the Planck scale $M_{\rm Pl}\sim 10^{18}~{\rm GeV}$, the electroweak scale $M_{\rm EW}\sim 10^{3}~{\rm GeV}$ and the vacuum energy scale $M_{\rm vac}\sim 10^{-12}~{\rm GeV}$.
@@ V.1.3 Hierarchy Puzzles: Seesaw Mechanism
    Here $M_{4}$ is the observed (4D) Planck mass $M_{\rm Pl}$, but what should we take for the higher-dimensional Planck mass $M_{4+n}$? Again following e.g. [4], we imagine that the enormous hierarchy between the $M_{\rm Pl}$ and $M_{\rm EW}$ is a manifestation of the fact that we are confined to a lower-dimensional slice of spacetime, rather than perceiving the higher-dimensional spacetime; in other words, whereas the lower-dimensional Planck mass has an enormous hierarchy relative to the electroweak scale ($M_{4}=M_{\rm Pl}\gg M_{\rm EW}$), the higher-dimensional Planck mass has no such hierachy ($M_{4+n}\approx M_{\rm EW}$). With this identification, Eq. (64) becomes
@@ Acknowledgements.
We thank Nikolay Bogachev, Mikhail Belolipetsky, Richard Borcherds, Gerald Höhn, Jury Radkovski, and Kendrick Smith for helpful discussions. SM would like to thank the School of Physics and Astronomy and the Higgs Center for Theoretical Physics at the University of Edinburgh for the hospitality in the following occasions, Fall 2023, July 2024, Fall 2024, July 2025, Fall 2025, without which finishing this work would not have be possible. He would also like to thank the Savvas Chamberlain Family Foundation and the Hellenic Heritage Foundation for their continuous support through the Anaximandros Fellowship (and personally Savvas and Christine Chamberlain, Maria Harper, Tula Alexopoulos, Michael Lagopoulos and Tony Lee) and the AHEPA Foundation for the 2024/25 Mavroyiannis Scholarship. LB is supported by the STFC Consolidated Grant “Particle Physics at the Higgs Centre.” Research at Perimeter Institute is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Economic Development and Innovation.
@@ Appendix A Reflection Symmetries of Odd Self-Dual Lorentzian Lattices $I_{s,1}$
We follow Vinberg’s approach in Refs. [68, 44] with the difference that our fundamental roots are denoted by $\mathbf{r}_{i}$ (instead of $\mathbf{e}_{j}$), while our standard orthonormal (Cartesian) basis vectors are denoted by ${\bf e}_{i}$ (instead of $u_{i}$); and, to align with the physicists’ notation, we label the Cartesian basis in $\mathbb{R}^{s,1}$ as $\{\mathbf{e}_{j}\}$ ($j=0,\dots,s$), with $\mathbf{e}_{0}$ corresponding to the temporal basis vector.
@@
    Figure 17: Coxeter-Dynkin diagrams for $\textrm{I}_{18,1}$ and $\textrm{I}_{19,1}$.
@@
    Now, we obtain the full set of fundamental roots ($\Sigma_{s}$) by adding ‘extra’ roots to the skeleton set ($\Sigma_{s}^{*}$) in such a way that the corresponding Coxeter diagram’s ($S_{4}$ or $S_{5}$) symmetry is maintained. In particular, the extra roots are the roots we previously excluded from $\Sigma_{s}^{0}$ to obtain $\Sigma_{s}^{\ast}$, as well as any additional roots needed to maintain the symmetry ($S_{4}$ or $S_{5}$, respectively) of the Coxeter-Dynkin diagram. In particular, all the extra vertices are either vertices of the 1st kind, related by ($S_{4}$ or $S_{5}$) symmetry to the root $\mathbf{r}_{s}$, or vertices of the 2nd kind, related by symmetry to the root $\mathbf{r}_{s+5}$.

### v2

@@ V.1.2 Hierarchy Puzzles: Cartoon Picture
Now let us see how this picture might be helpful in explaining some of the large hierarchies between three fundamental dimensional scales in Nature: the Planck scale $M_{\rm Pl}\sim 10^{18}~{\rm GeV}$, the electroweak scale $M_{\rm EW}\sim 10^{3}~{\rm GeV}$ and the vacuum energy scale $M_{\rm vac}\sim 10^{-12}~{\rm GeV}$.
@@ V.1.3 Hierarchy Puzzles: Seesaw Mechanism
    Here $M_{4}$ is the observed (4D) Planck mass $M_{\rm Pl}$, but what should we take for the higher-dimensional Planck mass $M_{4+n}$? Again following e.g. [4], we imagine that the enormous hierarchy between the $M_{\rm Pl}$ and $M_{\rm EW}$ is a manifestation of the fact that we are confined to a lower-dimensional slice of spacetime, rather than perceiving the higher-dimensional spacetime. In other words, whereas the lower-dimensional Planck mass has an enormous hierarchy relative to the electroweak scale ($M_{4}=M_{\rm Pl}\gg M_{\rm EW}$), the higher-dimensional Planck mass has no such hierarchy ($M_{4+n}\approx M_{\rm EW}$). With this identification, Eq. (64) becomes
@@ V.2 Other Speculations and Suggestions
  • Scale-symmetric quasicrystals, in particular the Ammann-Beenker tilings, have recently shown to be efficient in solving the general-case NP-complete problem of finding Hamiltonian cycles (i.e. closed loops on a graph visiting each vertex precisely once) [64]. Since many interesting problems in physics fall in the NP complexity class, this result hints towards the fact that quasicrystals are essentially as powerful (if not more) as periodic lattices in simplifying certain physical problems. It would be interesting to understand the graph-theoretic properties of spacetime quasicrystals and see what kind of problems their special properties can help us solve.
@@ Acknowledgements.
We thank Nikolay Bogachev, Mikhail Belolipetsky, Richard Borcherds, Gerald Höhn, Tobias Hartnick, Henna Koivusalo, Felix Flicker, Jury Radkovski, Ben Webster, and Kendrick Smith for helpful discussions. SM would like to thank the School of Physics and Astronomy and the Higgs Center for Theoretical Physics at the University of Edinburgh for the hospitality in the following occasions (Fall 2023, July 2024, Fall 2024, July 2025, Fall 2025), without which finishing this work would have been impossible. He would also like to thank the Savvas Chamberlain Family Foundation and the Hellenic Heritage Foundation for their continuous support through the Anaximandros Fellowship (and personally Savvas and Christine Chamberlain, Maria Harper, Tula Alexopoulos, Michael Lagopoulos and Tony Lee) and the AHEPA Educational Foundation for the 2024/25 Mavroyiannis Scholarship. LB is supported by the STFC Consolidated Grant “Particle Physics at the Higgs Centre.” Research at Perimeter Institute is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Economic Development and Innovation.
@@ Appendix A Reflection Symmetries of Odd Self-Dual Lorentzian Lattices $I_{s,1}$
We follow Vinberg’s approach in Refs. [70, 45] with the difference that our fundamental roots are denoted by $\mathbf{r}_{i}$ (instead of $\mathbf{e}_{j}$), while our standard orthonormal (Cartesian) basis vectors are denoted by ${\bf e}_{i}$ (instead of ${\bf u}_{i}$); and, to align with the physicists’ notation, we label the Cartesian basis in $\mathbb{R}^{s,1}$ as $\{\mathbf{e}_{j}\}$ ($j=0,\dots,s$), with $\mathbf{e}_{0}$ corresponding to the temporal basis vector.
@@
    Figure 17: Coxeter-Dynkin diagrams for (the skeletons of) $\textrm{I}_{18,1}$ and $\textrm{I}_{19,1}$ (see text for more information).
@@
    Now, we obtain the full set of fundamental roots ($\Sigma_{s}$) by adding ‘extra’ roots to the skeleton set ($\Sigma_{s}^{*}$) in such a way that the corresponding Coxeter diagram’s ($S_{4}$ or $S_{5}$) symmetry is maintained. In particular, the extra roots are the roots we previously excluded from $\Sigma_{s}^{(0)}$ to obtain $\Sigma_{s}^{\ast}$, as well as any additional roots needed to maintain the symmetry ($S_{4}$ or $S_{5}$, respectively) of the Coxeter-Dynkin diagram. In particular, all the extra vertices are either vertices of the 1st kind, related by ($S_{4}$ or $S_{5}$) symmetry to the root $\mathbf{r}_{s}$, or vertices of the 2nd kind, related by symmetry to the root $\mathbf{r}_{s+5}$.